for a normal population with a nean of u =80 and a standard devuatiin of 10, what is the probabilitu of obtaining a sample mean greater than M= 75 for a sample of n= 25 scores?
Answer #1

Let us consider the random variable ΧN N(μ = 80, σ = 10) , then its sample mean Xe Nu = 80, Vn= 10 V25 = 2) , where n=25

We know that the standard normal variate Z~N(0,1)

X-u P(X>75) = PO 75 - 80 -) = P(Z > -2.5) = 1- P(Z < -2.5)

1 - P(Z < -2.5) = 1- (1 - P(Z < 2.5)) = P(Z < 2.5) = 0.9938.

Therefore, the probability of obtaining a sample mean greater than M=75 is 0.9938.